Ann has some dimes and quarters. if she has 25 coins worth 4.45 how many of each type does she have?

Answer by Alan3354(30983)

Ann has some dimes and quarters. if she has 25 coins worth 4.45 how many of each type does she have?
----------------------------
If she had 25 dimes, she'd have $2.50
Each dime replace by a quarter adds 15 cents
445 - 250 = 195
195/15 = 13 quarters
10 dimes
-------------
The hard way:
d + q = 25
10d + 25q = 445
etc
Same answer